LAB VIL. PRELAB EXERCISE Name Section Consider the reaction, SO+32so+ Is. Data obtained in measuring rate of formation of Is are listed in the table. II1, M Initial rate, Ms 0.035 0.070 0.070 0.055 0.055 0.110 15 x 10 3.0 x 10 6.0 x 10 I. Determine the order of reaction with respect to S:Os 2What is the onder of reactin with reset t 3. Give the overall order of reaction. 4. Write the rate law for the reaction. 5. Determine the value for the rate constant, k. 6. What is the initial rate of reaction if IS:0s1 0.030M and I11-0.050 M? 7. When the rate of formation of ls is 3.0 x 105 M/s, what is the rate of disappearance of t? 37

I don't understand how to do these. Please help, thank you!
Answer #1

I am solving the first four parts as per Chegg guidelines, post multiple question to get the remaining answers

1)

Let the rate of the reaction is given by

Rate = k[S_2O_8^{2-}]^a [I^-]^b

We can write the rates of first and second experiment

1.5 * 10^{-5}= k[0.035]^a [0.055]^b ----(i)

3.0 * 10^{-5}= k[0.070]^a [0.055]^b ----(ii)

Dividing the first reaction by second we get

2^a = 2, a =1

Hence the order of the reaction wrt S2O8(2-) is equal to 1

2)

Let the rate of the reaction is given by

Rate = k[S_2O_8^{2-}]^a [I^-]^b

We can write the rates of first and third experiment

1.5 * 10^{-5}= k[0.035]^a [0.055]^b ----(i)

6.0 * 10^{-5}= k[0.070]^a [0.110]^b ----(ii)

Dividing the first reaction by second we get

2^a * 2^b = 4, b=1, since a is already calculated in the last part as 1

Hence the order of the reaction wrt I- is equal to 1

3)

Overall order of the reaction = Order wrt S2O8(2-) + Order wrt (I-) = 1 + 1 = 2

4)

Rate = k[S_2O_8^{2-}] [I^-]